The sum of the digits of a two-digir number is 12. If the digits are reversed, the new number is 18 more than the original number. Find the original number.

Answer by lwsshak3(11628)

The sum of the digits of a two-digir number is 12. If the digits are reversed, the new number is 18 more than the original number. Find the original number.
**
let u=units digit
let t=tens digit
t+u=12
t=12-u
10t+u=original number
10u+t=new number
original number+18=new number
10t+u+18=10u+t
10(12-u)+u+18=10u+(12-u)
120-10u+u+18=10u+12-u
18u=126
u=7
t=12-7=5
..
original number=57